A projectile is fired in such a way that its horizontal range is equal to its maximum height. What is the angle of projection?


1
Expert's answer
2021-11-05T18:20:04-0400

Equate them to figure out:


"R=H,\\\space\\
\frac{v^2\sin(2\theta)}{g}=\frac{v^2\sin\theta}{2g},\\\space\\
4\sin\theta\cos\theta=\sin\theta,\\\space\\
\theta=\arccos\frac 14=75.5°."